Find the range of an arrow that leaves the box at 50 m/s at ang angle of 50o above the horizontal?



1
Expert's answer
2021-10-08T09:21:59-0400

Given:

"v_0=50\:\rm m/s"

"\theta=50^{\circ}"


The range is given by

"R=\frac{v_0^2\sin2\theta}{g}=\frac{50^2\sin100^{\circ}}{9.8}=251\:\rm m"